From a lot of 10 missiles, 4 are selected at random and fired. If a lot contains 3 defective missiles that will not fire, what is the probability that :

i) All 4 will fire.

ii) At most 2 will not fire.

iii) How many defective missiles might we expect to be included among the 4 that are selected.


1
Expert's answer
2020-07-09T20:18:29-0400

i) The number of such chooses is the number of chooses of "4" missles from the "7" good ones, which equals "\binom{7}{4}". The total number of different chooses is "\binom{10}{4}", so the probability that all chosen missles will shoot equals "\binom{7}{4}/\binom{10}{4}=1/6" .

ii) "P(\text{at most }2 \text{ will not fire})=1-P(3\text{ missles will not fire})". The number of chooses of "4" missles such that there are "3" defective ones between them is "7" ("3" defected + an arbitrary missle from the left), so "P(\text{at most }2 \text{ will not fire})=1-7/\binom{10}{4}=29/30".

iii) Number the defective missles, and let "X_i=1" if the "i"-th missle is chosen and "X_i=0" if not. Then, by linearity of expectation, the expected number of defective missles chosen equals

"E(X_1+X_2+X_3)=E(X_1)+E(X_2)+E(X_3)=P(X_1\text{ is chosen})+P(X_2\text{ is chosen})+P(X_3\text{ is chosen})."

The number of choises such that the "i"-th defective missle is chosen is "\binom{9}{4}" (we choose this missle and "3" arbitrary from the "9" missles left), so

"E(X_1+X_2+X_3)=3\cdot\binom{9}{4}/\binom{10}{4}=6/5."